# Probing soft signals of gravitational-wave memory with space-based interferometers

### 30 Mar 2026 v1

@@ III.3 Displacement memory from quasi-circular non-precessing BBH merger
$h_{l,m}^{\text{mem}}(u)\approx\frac{16\pi}{R}\sqrt{\frac{(l-2)!}{(l+2)!}}\int^{u}dt^{\prime}\int d\Omega_{2}\,\frac{dE_{\text{gw}}}{dt^{\prime}\,d\Omega_{2}}\,Y_{lm}^{*}.$ (33)
Similar result also applies to other forms of null radiation. The angular energy flux distribution $\frac{dE_{\text{gw}}}{dt\,d\Omega_{2}}$ of the primary GW is determined by its energy density $\rho_{\text{gw}}=\left\langle\frac{1}{16\pi}|\dot{h}|^{2}\right\rangle$ in the wave zone, which can be written as a summation over $g=\{l_{1},m_{2},l_{2},m_{2}\}$ as [30]
@@
where the angle brackets denote an average over several GW wavelengths. For a quasi-periodic GW source, the null memory thus grows linearly in time. As discussed in Sec. II.1, the detectability of such a waveform is poor. For the search of null displacement memory, one is therefore interested in the system capable of producing a strong and predictable GW burst, but with negligible ordinary displacement memory. The ideal systems to observe are the binary black hole (BBH) mergers. In this section, we discuss the potentially observable soft waveforms of such systems. Our focus is not on identifying their nonlinear nature (which is impossible without observing the primary GW), but rather on some features of their low-frequency spectra. We consider quasi-circular non-precessing BBH systems and employ the numerical relativity surrogate model NRHybSur3dq8_CCE [96] (constructed for the mass ratio $q=m_{2}/m_{1}\in[1,10]$) to compute their gravitational waveforms.
@@
Long before the merger, 0PN approximation to the waveform is adequate. In this regime, the orbital evolution is given by $\dot{x}=Jx^{5}$, where $J\equiv 64\nu/(5M)$, the PN parameter $x\equiv(M\Omega)^{2/3}$, with $\Omega$ being the orbital frequency. Meanwhile, the gravitational radiation in a source frame with $Z$-axis aligned with the orbital angular momentum is dominated by the $h_{2,\pm 2}$ modes, corresponding to the quadrupole waveform. Due to the adiabaticity of orbital evolution, the memory mode after its steady accumulation also appears at 0PN order, and is dominated by the $h_{2,0}$ mode. This leads to the LO memory-mode waveform [92] with $h_{l,m}^{\text{(mem)}}(x=0)=0$: $h_{+}^{\text{(mem)}}(t)=Kx$, $h_{\times}^{\text{(mem)}}(t)=0$, where $K\equiv{\nu Ms^{2}_{\Theta}\,(17+c^{2}_{\Theta})}/(48R)$. At $t\geq t_{0}$, we are only interested in the change of $h_{+}^{\text{(mem)}}$ with respect to its initial value at $t_{0}$, which is given by $\Delta h_{+}^{\text{(mem)}}(t)\equiv h_{+}^{\text{(mem)}}(t)-h_{+}^{\text{(mem)}}(t_{0})=Kx_{0}\left[1-(1+4Jx_{0}^{4}\Delta t)^{-1/4}\right]$, where $x_{0}=x(t=t_{0})$ and $\Delta t=t-t_{0}$. A concrete example is presented in Fig. 6. As shown in the bottom panel, the initial memory-mode waveform is well-described by this 0PN approximation, which in turn can be approximated by a linear growth (expected for a quasi-periodic GW source) initially. The Newtonian regime of inspiral is followed by the late inspiral, merger, and ringdown, during which the memory accumulates much faster than the 0PN prediction before finally settling to a saturated value. Throughout this process, $h_{2,0}$ is the dominating memory mode. The top panel of Fig. 6 compares the memory-mode waveform with the full (memory + oscillatory) waveform. Although the oscillatory waveform is the absolutely dominating component before the final merger, the memory-mode waveform is the only final remenant of this event, if the observation starts before the merger. This is manifested in the frequency spectrum of the full waveform shown in Fig. 7, whose low-frequency part is contributed solely by the memory modes. In contrast, the low-frequency spectrum of $h_{\times}(t)$, dominated by the spin memory mode $h_{30}$ [69], is much weaker than that of $h_{+}(t)$. In practice, the low-frequency signal from spin memory would therefore be masked by that of displacement memory.
@@
As shown in the upper figure inset in Fig. 7, $\text{Im}\,\tilde{h}_{+}$ of the memory-mode waveform is almost purely imaginary until the frequency is sufficiently high, at which the spectrum has been overwhelmed by the oscillatory component that appears after $f\gtrsim\Omega(t_{0})/\pi$. Thus, a soft waveform with a negative real correction factor $C(f)$ provides a good approximation to $\tilde{h}(f)$ at low frequencies.
@@
Figure 6: Gravitational waveform of a quasi-circular BBH merger from NRHybSur3dq8_CCE, with $q=1$, $M=100\,M_{\odot}$, $\Theta=\pi/3$ and $R=400\,\text{Mpc}$ (the effect of cosmic expansion is not included). The top panel compares the full waveform (orange) with the memory-mode waveform due to $h_{2,0}$ and $h_{40}$ modes (purple), and the 0PN approximation to the oscillatory waveform (black dashed), for a non-spinning system. The black dot-dashed line shows the memory-mode waveform for the case of aligned dimensionless spins $\chi_{1,2}=0.8$. The blue dot-dashed line shows the 0PN approximation to the memory-mode waveform for an elliptical orbit [32] with eccentricity $e(t_{0})=0.05$ and semi-major axis $a(t_{0})=M/x_{0}$. The lower panel compares the $h_{2,0}$ mode (blue dashed) with its 0PN approximation (red dashed) and linear approximation (black dashed), for a non-spinning system.
@@
Figure 7: Frequency spectra of the waveforms in Fig. 6. The horizontal red (black) dashed line represents the soft waveform corresponding to $\Delta h_{+}$ of the full (reduced) time-domain waveform. The black solid line is the spectrum of the shortened waveform. We also show a local fit to the spectrum of the full waveform by a two-parameter logarithmic model $A+B\ln f$ (blue dashed line), the memory-mode spectrum for $q=8$ multiplied by 6.28 (green dot-dashed line), and the memory-mode spectrum for $q=1$ and aligned spins $\chi_{1,2}=0.8$ multiplied by 0.49 (black dot-dashed line). The blue dot-dashed line shows the 0PN oscillatory-mode spectrum ($f|\tilde{h}_{+}|\propto f^{-1/6}$) in the stationary phase approximation. The lower figure inset shows a zoomed-in view of the low-frequency region. The upper figure inset compares $f\,|\tilde{h}_{+}(f)|$ with $f\,|\text{Im}\,\tilde{h}_{+}(f)|$ (red dot-dashed line) of the memory-mode waveform.
It should be stressed that in this example the low-frequency spectrum does not faithfully reflect the memory amplitude, which is precisely due to the slow accumulation of memory during the early inspiral described above. The horizontal red dashed line in Fig. 7 indicates the value $|\Delta h_{+}^{\text{(mem)}}(t_{\text{final}})|$, while the black dashed line indicates the value $|\Delta h_{+}^{\text{(mem)}}(t_{\text{final}})-\Delta h_{+}^{\text{(mem)}}(t_{\text{mid}})|$, which is the memory height of the shortened waveform: $h_{+}^{\prime}(t)=[h_{+}(t)-h_{+}(t_{\text{mid}})]\,\Theta(t-t_{\text{mid}})$ $t_{\text{mid}}=t_{\text{merger}}-300\,\text{s}$ chosen so that $h_{2,0}$ can be described by the 0PN approximation for $t<t_{\text{mid}}$. We verified that the numerical spectrum of the waveform $h_{+}(t)-h_{+}^{\prime}(t)$, describing the growth prior to $t_{\text{mid}}$, agrees well with the analytical spectrum derived from the LO memory waveform, validating the accuracy of the FFT]. Neither horizontal line precisely matches the low-frequency limit of its corresponding spectrum, since $10^{-4}\,\text{Hz}$ is still finite. Even if $t_{\text{merger}}-t_{0}$ is fixed but we look at sufficiently low frequencies (which can be achieved by enlarging $t_{\text{final}}-t_{\text{merger}}$), we see a soft waveform with the correct height $|\Delta h_{+}^{\text{(mem)}}(t_{\text{final}})|$, which roughly becomes $h_{+}^{\text{(mem)}}(t_{\text{final}})-h_{+}^{\text{(mem)}}(t_{\text{mid}})+Kx(t_{\text{mid}})$ for $t_{0}\to$-$\infty$.
A small difference at the lowest frequencies is, however, the only place where the early evolution manifests itself, as the spectra of the shortened and full waveforms are barely distinguishable for $f>10^{-3}\,\text{Hz}$. Near the lowest frequencies, the spectrum can be approximated by a linear fit, whereas at moderately high frequencies it can be reasonably described by a logarithmic fit (which is much smaller than the linear correction observed at low frequencies), as shown by the blue dashed line in Fig. 7. For a given total mass $M$, as the mass ratio $q>1$ increases, the effective memory height decreases, while the relative correction $|C(f)|$ increases. This is illustrated in Fig. 7 by comparing the normalized spectra of systems with $q=1$ and $q=8$. Figs. 6-7 also illustrate that the total memory offset is insensitive to the orbital eccentricity in the early stage [32], while it can be significantly affected by the BH spins.
Enlarging $t_{\text{final}}-t_{\text{merger}}$ does not change the memory-mode spectrum, while enlarging $t_{\text{merger}}-t_{0}$ only leads to a small difference at the lowest frequencies if $x_{0}$ is small (as can be seen by comparing the shortened and full waveforms in Fig. 7). The impact of observation time on the memory-mode spectrum is thus small, provided that most of the memory is accumulated during the observational time interval. On the other hand, the lowest frequency covered by the oscillatory-mode spectrum is $f\approx\Omega(t_{0})/\pi$. In the 0PN regime, we have ${\Omega(t_{0}-\Delta t)}/{\Omega(t_{0})}=\ell$, with
@@
Assuming that $\Omega(t_{0})/\pi=1\,\text{Hz}$, for $M=100\,M_{\odot}$ and $\nu=q/(1+q)^{2}=1/4$, $\ell(\Delta t=10^{-4}\,\text{yr})\approx 0.62$, $\ell(\Delta t=10^{-3}\,\text{yr})\approx 0.3$, while $\ell(\Delta t=1\,\text{yr})\approx 0.02$. For smaller $M$, $\Omega(t_{0})/\pi=1\,\text{Hz}$ corresponds to an earlier stage of the evolution. Therefore for the quasi-circular stellar-mass BBH mergers and $\Delta t<10^{-3}\,\text{yr}$, typically $\Omega(t_{0}-\Delta t)/\pi\sim 1\,\text{Hz}$.
Due to the mass-scaling symmetry of the orbital dynamics and gravitational waveforms for fixed dimensionless spins and mass ratio, the spectra of observed memory-mode waveforms of sufficiently long duration in an expanding Universe (see Sec. III.1) exhibit the usual mass–redshift degeneracy:
$\tilde{h}^{\prime}(f,M^{\prime},d_{L})=\left[\frac{(1+z)M^{\prime}}{M}\right]^{2}\tilde{h}\left(\frac{(1+z)M^{\prime}}{M}f,M,d_{L}\right).$

### 13 Apr 2026 v2

@@ III.3 Displacement memory from BBH merger
$h_{l,m}^{\text{(mem)}}(u)\approx\frac{16\pi}{R}\sqrt{\frac{(l-2)!}{(l+2)!}}\int^{u}dt^{\prime}\int d\Omega_{2}\,\frac{dE_{\text{gw}}}{dt^{\prime}\,d\Omega_{2}}\,Y_{lm}^{*},$ (27)
where $d\Omega_{2}(\theta,\phi)\equiv\sin\theta\,d\theta\,d\phi$. A similar result holds for other forms of null radiation, but the gravitational case is special since it arises from the self-interaction of GWs. The angular distribution of the energy flux $\frac{dE_{\text{gw}}}{dt\,d\Omega_{2}}$ of the primary GW is determined by its energy density $\rho_{\text{gw}}=\left\langle\frac{1}{16\pi}|\dot{h}|^{2}\right\rangle$ in the wave zone, which can be written as a summation over $g=\{l_{1},m_{1},l_{2},m_{2}\}$ as [33]
@@
where the angle brackets denote an average over several GW wavelengths. For a quasi-periodic GW source, the null memory thus grows linearly in time. As discussed in Sec. II.1, the detectability of such a waveform is poor. For the search of null displacement memory, one is therefore interested in the system capable of producing a strong and predictable GW burst, but with negligible ordinary displacement memory. The ideal systems to observe are the binary black hole mergers. In this section, we discuss some features of the potentially observable low-frequency spectra of such systems. We consider quasi-circular non-precessing BBH systems and employ the numerical relativity surrogate model NRHybSur3dq8_CCE [103] (constructed for the mass ratio $q=m_{2}/m_{1}\in[1,10]$) to compute their gravitational waveforms.
@@
Long before the merger, 0PN approximation to the waveform is adequate. In this regime, the orbital evolution is given by $\dot{x}=Jx^{5}$, where $J\equiv 64\nu/(5M)$, the PN parameter $x\equiv(M\Omega)^{2/3}$, with $\Omega$ being the orbital frequency. Meanwhile, the gravitational radiation in a source frame with $Z$-axis aligned with the orbital angular momentum is dominated by the $h_{2,\pm 2}$ modes, corresponding to the quadrupole waveform. Due to the adiabaticity of orbital evolution, the memory after its steady accumulation also appears at 0PN order, and is dominated by the $h_{2,0}$ mode. The LO memory-mode waveform with $h_{l,m}^{\text{(mem)}}(x=0)=0$ is given by [^8] For a primary GW given by $h_{2,\pm 2}(t)=Ae^{\mp 2i\Omega t}$ with $A\in\mathbb{R}$, Eq. (27) yields $\dot{h}_{\times}^{\text{(mem)}}=0$ and $\dot{h}_{+}^{\text{(mem)}}=RA^{2}\Omega^{2}s^{2}_{\Theta}\,(17+c^{2}_{\Theta})/(48\pi)$. For the 0PN oscillatory waveform of a circular binary, $A=8\sqrt{\pi/5}\,\nu Mx(t)/R$. Note that the oscillatory waveform of a vector gravitational atom in a saturated superradiant ground state [83, 15] is similar, with $A\approx A(t_{0})/[1+(t-t_{0})/\tau_{\text{gw}}]$, so that $\Delta h_{+}^{\text{(mem)}}\propto 1-\tau_{\text{gw}}/(t-t_{0}+\tau_{\text{gw}})$. [99] $h_{+}^{\text{(mem)}}(t)=Kx$ and $h_{\times}^{\text{(mem)}}(t)=0$, where $K\equiv{\nu Ms^{2}_{\Theta}\,(17+c^{2}_{\Theta})}/(48R)$. The change of $h_{+}^{\text{(mem)}}$ with respect to its initial value at $t_{0}$ is given by $\Delta h_{+}^{\text{(mem)}}(t)\equiv h_{+}^{\text{(mem)}}(t)-h_{+}^{\text{(mem)}}(t_{0})=Kx_{0}\left\{1-[1+4Jx_{0}^{4}(t-t_{0})]^{-1/4}\right\}$, where $x_{0}=x(t=t_{0})$. A concrete example with $M=100\,M_{\odot}$ and $q=1$ is shown in Fig. 6. As shown in the bottom panel, the initial memory-mode waveform is well-described by this 0PN approximation, which in turn can be approximated by a linear growth initially. The Newtonian regime of inspiral is followed by the late inspiral and merger, during which the memory accumulates much faster than the 0PN prediction, before finally settling to a saturated value in the ringdown phase. Throughout this process, $h_{2,0}$ is the dominating memory mode. The top panel of Fig. 6 compares the memory-mode waveform with the full (memory + oscillatory) waveform. Although the oscillatory waveform is the dominating component before the final merger, the memory-mode waveform is the only final remenant of this event, if the observation starts before the merger. This is manifested in the frequency spectrum of the full waveform shown in Fig. 7, whose low-frequency part is contributed solely by the memory modes. The low-frequency spectrum of $h_{\times}(t)$ is dominated by the spin memory mode $h_{30}$ [74], but it is much weaker than that of $h_{+}(t)$. In practice, the low-frequency signal from spin memory would therefore be masked by that of displacement memory.
@@
As shown in the upper figure inset in Fig. 7, $\tilde{h}_{+}$ of the memory-mode waveform is almost purely imaginary until the frequency is sufficiently high, at which the spectrum has been overwhelmed by the oscillatory component that appears after $f\gtrsim\Omega(t_{0})/\pi$. Thus, a soft waveform with a negative real correction factor $C(f)$ provides a good approximation to $\tilde{h}(f)$ at low frequencies.
@@
Figure 6: Gravitational waveform of a quasi-circular BBH merger from NRHybSur3dq8_CCE, with $q=1$, $M=100\,M_{\odot}$, $\Theta=\pi/3$ and $R=400\,\text{Mpc}$ (the effect of cosmic expansion is not included). The top panel compares the full waveform (orange) with the memory-mode waveform due to $h_{2,0}$ and $h_{4,0}$ modes (purple), and the 0PN approximation to the oscillatory waveform (black dashed), for a non-spinning system. The black dot-dashed line shows the memory-mode waveform for the case of aligned dimensionless spins $\chi_{1,2}=0.8$. The blue dot-dashed line shows the 0PN approximation to the memory-mode waveform for an elliptical orbit [35] with eccentricity $e(t_{0})=0.05$ and semi-major axis $a(t_{0})=M/x_{0}$. The lower panel compares the $h_{2,0}$ mode (blue dashed) with its 0PN approximation (red dashed) and linear approximation (black dashed), for a non-spinning system.
@@
Figure 7: Frequency spectra of the waveforms in Fig. 6. The horizontal red (black) dashed line corresponds to $\Delta h_{+}$ of the full (shortened) time-domain waveform in $t_{0},t_{\text{final}}]$; the horizontal gray dashed line corresponds to $(\Delta h_{+})_{\text{max}}$. The black solid line is the spectrum of the shortened waveform. We also show a local fit to the spectrum of the full waveform by a two-parameter logarithmic model $A+B\ln f$ (blue dashed line), the memory-mode spectrum for $q=8$ multiplied by 6.28 (green dot-dashed line), and the memory-mode spectrum for $q=1$ and aligned spins $\chi_{1,2}=0.8$ multiplied by 0.49 (black dot-dashed line). The blue dot-dashed line shows the 0PN oscillatory-mode spectrum ($f|\tilde{h}_{+}|\propto f^{-1/6}$) in the stationary phase approximation. The lower figure inset shows a zoomed-in view of the low-frequency region. The upper figure inset compares $f\,|\tilde{h}_{+}(f)|$ with $f\,|\text{Im}\,\tilde{h}_{+}(f)|$ (red dot-dashed line) of the memory-mode waveform. $f_{\text{ISCO}}\equiv\sqrt{M/(6M)^{3}}/\pi\approx 22\,\text{Hz}$, and the maximum oscillation frequency $\sim 100\,\text{Hz}$.
Due to the slow accumulation of memory during the early inspiral, the low-frequency spectrum in Fig. 7 does not faithfully reflect the memory amplitude. The horizontal red dashed line in Fig. 7 indicates the value $|\Delta h_{+}^{\text{(mem)}}(t_{\text{final}})|/(2\pi)$, while the black dashed line indicates the value $|\Delta h_{+}^{\text{(mem)}}(t_{\text{final}})-\Delta h_{+}^{\text{(mem)}}(t_{\text{mid}})|/(2\pi)$, which is the memory amplitude of the shortened waveform (with zero padding for $t<t_{\text{mid}}$): $h_{+}^{\prime}(t)=[h_{+}(t)-h_{+}(t_{\text{mid}})]\,\Theta(t-t_{\text{mid}})$, with [^9] This is chosen so that $h_{2,0}$ can be described by the 0PN approximation for $t<t_{\text{mid}}$. We verified that the numerical spectrum of the waveform $h_{+}(t)-h_{+}^{\prime}(t)$, describing the growth prior to $t_{\text{mid}}$, agrees well with the analytical spectrum derived from the 0PN memory-mode waveform, validating the accuracy of the DFT. $t_{\text{mid}}=t_{\text{merger}}-300\,\text{s}$. Neither horizontal line precisely matches the low-frequency limit of its corresponding spectrum, although the differences are small. A small difference between the spectra of the shortened and full waveforms appears only below $f=10^{-2}\,\text{Hz}$, reflecting the early evolution during $t_{0},t_{\text{mid}}]$. We can approximately construct $\tilde{h}^{\text{(mem)}}_{\infty}(f)$ using $h_{+}^{\text{(mem)}}(t)$ in $t_{\text{mid}},t_{\text{final}}]$ and the analytical spectrum of the 0PN memory-mode waveform, which is almost indistinguishable from the blue dashed line. At sufficiently low frequencies, $\tilde{h}^{\text{(mem)}}_{\infty}(f)$ approaches the soft waveform corresponding to $(\Delta h_{+})_{\text{max}}\equiv h_{+}^{\text{(mem)}}(\infty)-h_{+}^{\text{(mem)}}(-\infty)\approx h_{+}^{\text{(mem)}}(t_{\text{final}})-h_{+}^{\text{(mem)}}(t_{\text{mid}})+Kx(t_{\text{mid}})$, see the upper figure inset in Fig. 7. Since the memory growth rate during the ringdown phase is exponentially damped, the exact LO correction factor $\mathcal{C}(f)$ (for $t_{0}=$-$\infty$ and $t_{\text{final}}=\infty$) is dominated by the early-time growth and scale as $f^{1/4}$. At $f\gtrsim f_{\text{min}}\sim 1/(t_{\text{final}}-t_{0})$, the spectrum can be approximated by a linear fit, whereas at moderately high frequencies it can be roughly described by a logarithmic fit, as shown by the blue dashed line in Fig. 7. Polynomial fits appear to be unsuitable for describing the spectrum in the frequency band of interest. For a given total mass $M$, as the mass ratio $q>1$ increases, $\Delta h_{+}^{\text{(mem)}}(t_{\text{final}})$ decreases, while the relative correction $|C(f)|$ increases. This is illustrated in Fig. 7 by comparing the normalized spectra of systems with $q=1$ and $q=8$. Figs. 6-7 also illustrate that the 0PN memory growth after $t_{0}$ is insensitive to a small initial orbital eccentricity [35], while the later growth can be significantly affected by the BH spins. As the observation time increases, lower frequency components of the memory mode can be resolved, while the lowest frequency of the oscillatory mode is $f\approx\Omega(t_{0})/\pi$. In the 0PN regime, we have ${\Omega(t_{0}-\Delta t)}/{\Omega(t_{0})}=\ell$, with
@@
Assuming that $\Omega(t_{0})/\pi=1\,\text{Hz}$, for $M=100\,M_{\odot}$ and $\nu=1/4$, $\ell(\Delta t=10^{-4}\,\text{yr})\approx 0.62$, $\ell(\Delta t=10^{-3}\,\text{yr})\approx 0.3$, while $\ell(\Delta t=1\,\text{yr})\approx 0.02$. For smaller $M$, $\Omega(t_{0})/\pi=1\,\text{Hz}$ corresponds to an earlier stage of the evolution. Therefore for the quasi-circular stellar-mass BBH mergers and $\Delta t<10^{-3}\,\text{yr}$, typically $\Omega(t_{0}-\Delta t)/\pi\sim 1\,\text{Hz}$.
Due to the mass-scaling symmetry of the orbital dynamics and gravitational waveforms for fixed dimensionless spins and mass ratio, the memory-mode spectra in an expanding universe (see Sec. III.1) exhibit the usual mass–redshift degeneracy:
$\tilde{h}_{\infty}^{\prime}(f,M^{\prime},d_{L})=\left[\frac{(1+z)M^{\prime}}{M}\right]^{2}\tilde{h}_{\infty}\left(\frac{(1+z)M^{\prime}}{M}f,M,d_{L}\right).$
Neglecting cosmic expansion, for $(q,M/M_{\odot},d_{L}/\text{pc})=(1,10^{2},4\times 10^{8})$, $(\Delta h_{+})_{\text{max}}\sim 3\times 10^{-22}$. Subsolar-mass compact binary mergers (e.g., of primordial BHs) may occur at much closer distances and thus be detectable [30]. E.g., $(q,M/M_{\odot},d_{L}/\text{pc})=(1,2,10^{4})$ results in $(\Delta h_{+})_{\text{max}}\sim 2\times 10^{-19}$, while $(q,M/M_{\odot},d_{L}/\text{pc})=(1,4\times 10^{-4},10^{-2})$ [46] results in $(\Delta h_{+})_{\text{max}}\sim 5\times 10^{-17}$. In the latter case, the memory-mode spectrum below 0.1 Hz may closely resemble a soft waveform.